If the actual budget deficit is $100 billion, the economy is operating $250 billion below its potential, and the marginal tax rate is 20 percent, then:

a. There is a passive surplus of $150 billion, and a structural deficit of $50 billion
b. There is a passive surplus of $50 billion, and a structural deficit of $150 billion
c. There is a passive deficit of $50 billion, and a structural deficit of $50 billion
d. There is a passive deficit of $50 billion, and a structural deficit of $150 billion
1
Expert's answer
2012-05-09T08:50:03-0400
d. There is a passive deficit of $50 billion, and a structural deficit of $150 billion
If the actual budget deficit is $100 billion, the economy is operating $250 billion below its
potential, and the marginal tax rate is 20 percent, then passive deficit = 100 -
250*0.2 = 50, structural deficit = 250 - 50 - 50 = 150